# ShortcutBreaker: Low-Rank Noisy Bottleneck with Global Perturbation Attention for Multi-Class Unsupervised Anomaly Detection

### v1

@@ Abstract
Multi-class unsupervised anomaly detection (MUAD) has garnered growing research interest, as it seeks to develop a unified model for anomaly detection across multiple classes—eliminating the need to train separate models for distinct objects and thereby saving substantial computational resources. Under the MUAD setting, while advanced Transformer-based architectures have brought significant performance improvements, identity shortcuts persist: they directly copy inputs to outputs, narrowing the gap in reconstruction errors between normal and abnormal cases, and thereby making the two harder to distinguish. Therefore, we propose ShortcutBreaker, a novel unified feature-reconstruction framework for MUAD tasks, featuring two key innovations to address the issue of shortcuts. First, drawing on matrix rank inequality, we design a low-rank noisy bottleneck (LRNB) to project high-dimensional features into a low-rank latent space, and theoretically demonstrate its capacity to prevent trivial identity reproduction. Second, leveraging ViT’s global modeling capability instead of merely focusing on local features, we incorporate a global perturbation attention to prevent information shortcuts in the decoders. Extensive experiments are performed on four widely used anomaly detection benchmarks, including three industrial datasets (MVTec-AD, ViSA, and Real-IAD) and one medical dataset (Universal Medical). The proposed method achieves a remarkable image-level AUROC of 99.8%, 98.9%, 90.6%, and 87.8% on these four datasets, respectively, consistently outperforming previous MUAD methods across different scenarios. Our code will be released.
@@ 1 Introduction
Before deep learning, anomaly detection relied on traditional techniques: density-based methods (Breunig et al., 2000; Guan et al., 2015), distance-based methods (Knorr et al., 2000; Angiulli et al., 2005), and statistics-based methods (Hido et al., 2011; Rousseeuw & Hubert, 2011). Current state-of-the-art methods employ deep learning networks pre-trained on ImageNet (Deng et al., 2009) to capture discriminative features. Feature reconstruction methods (Guo et al., 2023; 2024; Deng & Li, 2022) reconstruct encoder-extracted features, assuming accurate reconstruction of normal regions but failure on unseen anomalies. Memory matching methods (Yi & Yoon, 2020; Defard et al., 2021; Roth et al., 2022) memorize training-set normal features for inference matching, with those in (Defard et al., 2021; Roth et al., 2022) using pre-trained encoders for discriminative features. Pseudo-anomaly methods (Liu et al., 2023; Li et al., 2021; Zavrtanik et al., 2021) convert UAD to a supervised task by generating pseudo anomalies via noise addition to normal images/features. Hybrid methods (Tien et al., 2023; Zhao et al., 2023) integrate normalizing flows (Zhao et al., 2023) or pseudo noise (Tien et al., 2023) into feature reconstruction for UAD.
@@
Figure 2: The visual patterns in the medical field are richer than those in the industrial field. (a) Visualization results, (b) T-SNE plots (Van der Maaten & Hinton, 2008), and (c) distributions of feature distances for the Universal Medical and ViSA datasets (Zou et al., 2022) are presented. These experimental results help observe the diversity of each dataset. For T-SNE plots, we extract the final features from the pre-trained ResNet-50 (He et al., 2016). Moreover, we use LPIPs (Zhang et al., 2018) to compare feature distances between individual image pairs.
@@
Over the past three years, substantial progress has been made in multi-class unsupervised anomaly detection (MUAD), with explorations into pretrained vision transformers (ViTs) (Zhang et al., 2023; Guo et al., 2025), state space models (Mambas) (He et al., 2024a), diffusion models (Yin et al., 2023; He et al., 2024b), and other approaches (Guo et al., 2024; Zhao, 2023; He et al., 2024c). Nevertheless, the multi-class setting inevitably induces identity mapping in most methods, resulting in performance degradation. While previous studies (You et al., 2022; Guo et al., 2025) have endeavored to mitigate shortcut learning, the efficacy of their proposed techniques in complex and highly diverse scenarios remains limited (see Fig. 1 (c)-(d)). This is particularly evident in datasets such as Real-IAD and Universal Medical, where the abundance of normal patterns exacerbates the aforementioned issue (You et al., 2022). The effectiveness of simple noise operations (You et al., 2022; Guo et al., 2025) is diminished due to the enhanced noise robustness acquired through exposure to diverse visual patterns. Additionally, the neighbor mask attention mechanism (You et al., 2022) is specifically tailored for features locally extracted by CNNs, and is not suitable for the global modeling structure - Transformer. Consequently, to enhance performance, our work aims to efficiently address the problem of identity shortcuts. Furthermore, why are the Real-IAD and Universal Medical datasets more complex? We elaborate on this as follows: It is apparent that in comparison to widely adopted industrial datasets—MVTec-AD (encompassing 15 classes) (Bergmann et al., 2019) and ViSA (comprising 12 classes) (Zou et al., 2022)—Real-IAD (Wang et al., 2024) exhibits more diverse normal patterns, featuring 30 object categories and 5 camera views. In contrast to standardized industrial images, although the Universal Medical dataset (He et al., 2024b) consists of only 3 categories, the inherent heterogeneity within medical normal samples enriches the visual features (e.g., variations in organ size and shape across different demographics, as illustrated in Fig. 2(a)). Moreover, the Universal Medical dataset displays a sparser t-SNE plot and a wider dispersion of feature distances compared to ViSA (see Fig. 2 (b) and (c)), indicating a higher degree of intrinsic diversity within medical data.
@@
In this paper, we propose a simple yet effective framework for the MUAD task, named ShortcutBreaker, built on the advanced DINO-pretrained vision transformer with two key innovations. First, drawing on two observed properties of low-rank matrix decomposition, we design a low-rank noisy bottleneck (LRNB) to effectively mitigate the identity mapping issue. Within LRNB, the low-rank property theoretically circumvents shortcut learning of unseen patterns, while learnable matrix parameters are optimized to reconstruct normal patterns. Second, considering that ViT-extracted features share similar global information, we introduce a global perturbation attention (GPA) mechanism to curb information leakage from input to output in the decoder. In GPA, a global redistribution operation and a global-self-masking mechanism force the decoder to learn reconstruction based on longer-range and impaired features. To validate the effectiveness of the proposed method, we conduct extensive experiments on four publicly available datasets: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Universal Medical (He et al., 2024b), and Real-IAD (Wang et al., 2024). As presented in Fig. 1, our ShortcutBreaker achieves the highest image-level AUC: 99.8%, 98.9%, 88.2%, and 90.6% on these four datasets, respectively. Notably, on the complex Universal Medical and Real-IAD datasets, it outperforms previous methods by a significant margin. In summary, our contributions are:
@@
  • We propose a global perturbation attention mechanism, which effectively prevents shortcut learning in the decoder via global redistribution and global-self-masking operations.
@@
2 Related works
2.1 Single Class UAD
Recently, reconstruction mechanisms have dominated unsupervised anomaly detection (UAD) methods, detecting outliers by evaluating reconstruction errors between input and output pixels or features. These approaches include pixel reconstruction methods (Schlegl et al., 2017a; b) and feature reconstruction methods (Deng & Li, 2022; Guo et al., 2023; 2024). Pixel reconstruction methods aim to recreate normal images from scratch using a generative model framework. They assume that reconstructing normal samples will produce low reconstruction errors, while abnormal samples—differing from training data—will result in higher errors. Feature reconstruction methods suggest that latent features extracted from a pre-trained encoder provide more discriminative and robust representations compared to pixel-level reconstructions. Building on the success of such methods, many subsequent approaches (e.g., EDC (Guo et al., 2023) and AE-FLOW (Zhao et al., 2023)) leverage this foundation to enhance anomaly detection. For instance, EDC employs an unfrozen pre-trained encoder, which is fine-tuned to adapt to the target domain while using a stop-gradient operation to retain important domain-relevant features. AE-FLOW further extends this idea by combining feature map reconstruction errors with distribution distances (measured via normalizing flows) as the final anomaly score. However, simply applying single-class methods as part of this one-class-one-model scheme is memory-intensive (especially as the number of classes grows) and ill-suited for scenarios with high intra-class diversity in normal samples, such as multi-class UAD (You et al., 2022).
@@ 2.2 Multi Class UAD
UniAD (You et al., 2022) pioneers the field of multi-class unsupervised anomaly detection (MUAD), presenting a unified model capable of detecting anomalies across various classes. Most subsequent works have explored advanced modules to build better reconstruction models for MUAD. For example, LafitE (Yin et al., 2023) and DiAD (He et al., 2024b) further advance the MUAD task by leveraging the generative power of diffusion models to better capture anomalies across multiple classes. ViTAD (Zhang et al., 2023) and MambaAD (He et al., 2024a) develope feature reconstruction-based MUAD methods using recently advanced modules: Vision Transformer (Zhang et al., 2023) and State Space Model (He et al., 2024a), respectively. Few methods explicitly aim to address the identity mapping issue. UniAD (You et al., 2022) counteracts this through techniques like feature jittering and neighbor-masked attention. Dinomaly (Guo et al., 2025) employs components such as noisy bottleneck dropout to disrupt input feature replication. However, these noise operations have limited effectiveness in preventing shortcut learning when training on complex medical/industrial datasets (Universal Medical and Real-IAD), and neighbor-masked attention—specifically designed for CNN-extracted features—is not suitable for advanced ViT architectures (Guo et al., 2025; Zhang et al., 2023).
@@ 3.1 Overview
Following ViTAD and Dinomaly (Zhang et al., 2023; Guo et al., 2025), our ShortcutBreaker is a feature-reconstruction framework and constructed based on the vision transformer (ViT) structure. As depicted in Fig. 3(a), ShortcutBreaker consists of a pre-trained encoder, a bottleneck, and a decoder. The DINO-pretrained ViT model (Darcet et al., 2023) is utilized as the encoder, capturing informative feature tokens to facilitate subsequent reconstruction. The bottleneck is built upon a multi-layer perceptron (MLP), integrating multi-scale feature representations from intermediate encoder layers. The decoder maintains structural similarity with the encoder, employing transformer layers to reconstruct feature maps. As illustrated in Fig. 3(d), the encoder remains frozen during training, while the remaining components of ShortcutBreaker are optimized to reconstruct the encoder’s feature tokens by minimizing inter-token representation discrepancies. During inference, this configuration enables accurate reconstruction of normal patterns but exhibits reconstruction failures in anomalous regions. The anomaly localization map is generated by computing 1-CS (where CS denotes cosine similarity) between the original and reconstructed feature tokens and then reshaping the result. The anomaly detection score is subsequently determined by the maximum value within this anomaly map.
Figure 3: Flowchart of our ShortcutBreaker: (a) the overall structure of our proposed method; (b) the structure of the Low-Rank Noisy Bottleneck (LRNB); (c) the structure of the Global Perturbation Attention (GPA), which consists of GRD (global redistribution) and GSM (global-self masking); and (d) the training and inference pipelines.
@@ 3.2 Low-Rank Noisy Bottleneck (LRNB)
As discussed in the introduction, under the MUAD setting, identity shortcuts readily occur, narrowing the gap in anomaly scores between normal and abnormal cases. Previous methods (Guo et al., 2025; You et al., 2022) propose perturbing extracted feature tokens to compel the network to reconstruct from information-impaired normal features, rather than directly copying them. However, these noisy operations are often limited in effectively preventing shortcut learning in complex scenarios. We hypothesize that richer visual patterns inherent in the data enhance the reconstruction model’s robustness to noise. Consequently, we aim to design a novel paradigm for avoiding shortcuts that goes beyond solely relying on noisy operations. Through empirical observation of low-rank matrix decomposition in LoRA (Hu et al., 2022) (or the encoder-decoder structure in Auto-Encoder (He et al., 2022)), we identified two inherent properties particularly suitable for addressing the identity mapping issue.
Property 1 stems from the constrained low-rank latent space, which effectively circumvents identity shortcuts. In deep learning, nonlinear activations are generally used to improve performance, so we need to prove Property 1 under non-linear transformations, primarily using the Jacobian matrix (Goodfellow et al., 2016) for this proof.
To achieve identity mapping of $x\in\mathbb{R}^{d}$, we require that $g\circ f(x)=x$ holds for all $x\in\mathbb{R}^{n}$. Differentiating both sides with respect to xxyields the Jacobian equation:
@@
$r\left(J_{g}(f(x))\cdot J_{f}(x)\right)\leq\min\left(r(J_{g}(f(x))),r(J_{f}(x))\right)\leq k$} (3)
@@
Here, $r(\cdot)$ indicate the rank of the input. Under the low-rank adaptation module, we have $k<d$, Substituting into Eq. equation 3 implies:
$r\left(J_{g}(f(x))\cdot J_{f}(x)\right)\leq k<d$} (4)
@@
This contradicts Eqs. equation 2 that $det(I_{d\times d})$ requires rank=$d$. Consequently, no such functions $f$ and $g$ exist.
@@
Property 2 This property emerges from training exclusively on normal samples, during which the learnable $g$ and $f$ layers are optimized to specifically enhance the reconstruction of normal patterns. This optimization process minimizes adverse effects on the reconstruction fidelity of normal cases, thereby achieving a more favorable trade-off between the reconstruction of normal and abnormal samples.
Specifically, we adapt the LoRA framework (Hu et al., 2022) by implementing Multi-Layer Perceptrons (MLPs) with nonlinear activations to parameterize functions $g$ and $f$, constructing an auto-encoder-like structure. As illustrated in Fig. 3, the modules $g$ and $f$ exhibit symmetrical architectures, where $g$ comprises $i$ downsampling blocks (DBs) and $f$ contains $i$ upsampling blocks (UBs). Within DBs, the output token count of the $i$-th block is halved relative to the (i1)(i-1)-th block, while UBs correspondingly double the token count relative to their preceding block. Furthermore, we explore integrating feature-level noisy operations via Dropout  (Guo et al., 2025), and introduce additional perturbations for better performance.
3.3 Global Perturbation Attention (GPA)
Figure 4: Comparison of sigmoid and softmax Functions on a 2D attention map with Gaussian distribution, (a), (b), (c) are the 3D plot of original data, after softmax and sigmoid respectively, (d), (e), (f) are the corresponding heatmaps.
The self-attention mechanism in vanilla Transformer inherently facilitates identity mapping issues (You et al., 2022), as it permits unrestricted interaction between feature tokens and their own representations. Therefore, we propose a Global Perturbation Attention (GPA) in the decoder to prevent shortcut pathways in ViT architectures. As illustrated in Fig. 3, GPA comprises two core components: (1) global redistribution and (2) joint global-self masking.
Global Redistributing (GRD)
In the vanilla self-attention block (Eq. (5)), Softmax function is applied to the query-key attention map to conduct non-linear normalization (Alexey, 2020).
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (5)

### v2

# ShortcutBreaker: Low-Rank Noisy Bottleneck and Frequency Filtering Block for Multi-Class Unsupervised Anomaly Detection

@@ Abstract
Multi-class Unsupervised Anomaly Detection (MUAD) has attracted increasing research interest for its ability to develop a unified model across multiple classes, significantly reducing the computational cost of training separate models for distinct objects. However, despite the performance gains offered by advanced Transformer-based architectures, the identity shortcut issue persists. These models tend to directly copy inputs to outputs, narrowing the reconstruction error gap between normal and abnormal cases and hindering distinguishability. To address this, we propose ShortcutBreaker, a framework designed to disrupt such shortcuts through two core innovations. First, we introduce a Low-Rank Noisy Bottleneck (LRNB) derived from matrix rank inequality, which theoretically prevents trivial identity mapping by perturbing features across two distinct reconstruction groups. Second, we design a Frequency-Filter Unit (FFU) that integrates frequency spectral filtering and global-local masking. This mechanism further enlarges the anomaly score gap, with less computational cost than standard self-attention. Extensive evaluations on four benchmarks (MVTec-AD, ViSA, Real-IAD, and Universal Medical) demonstrate the superiority of ShortcutBreaker. Our code will be released.
Machine Learning, ICML
@@
Before deep learning, anomaly detection relied on traditional techniques: density-based methods (Breunig et al., 2000; Guan et al., 2015),distance-based methods (Knorr et al., 2000; Angiulli et al., 2005), and statistics-based methods (Hido et al., 2011; Rousseeuw & Hubert, 2011). Current state-of-the-art methods employ deep learning networks pre-trained on ImageNet (Deng et al., 2009) to capture discriminative features. Feature reconstruction methods (Guo et al., 2023, 2024; Deng & Li, 2022) reconstruct encoder-extracted features, assuming accurate reconstruction of normal regions but failure on unseen anomalies. Memory matching methods (Yi & Yoon, 2020; Defard et al., 2021; Roth et al., 2022) memorize training-set normal features for inference matching, with those in (Defard et al., 2021; Roth et al., 2022) using pre-trained encoders for discriminative features. Pseudo-anomaly methods (Liu et al., 2023; Li et al., 2021; Zavrtanik et al., 2021) convert UAD to a supervised task by generating pseudo anomalies via noise addition to normal images/features. Hybrid methods (Tien et al., 2023; Zhao et al., 2023) integrate normalizing flows (Zhao et al., 2023) or pseudo noise (Tien et al., 2023) into feature reconstruction for UAD.
@@
Figure 2: The visual patterns in the medical field are richer than those in the industrial field. (a) Visualization results (b) T-SNE plots (Van der Maaten & Hinton, 2008), and (c) distributions of feature distances for the Universal Medical and ViSA datasets (Zou et al., 2022) are presented. These experimental results help observe the diversity of each dataset. For T-SNE plots, we extract the final features from the pre-trained ResNet-50 (He et al., 2016). Moreover, we use LPIPs (Zhang et al., 2018) to compare feature distances between individual image pairs.
@@
Over the past three years, substantial progress has been made in multi-class unsupervised anomaly detection (MUAD), with explorations into pretrained vision transformers (ViTs) (Zhang et al., 2023; Guo et al., 2025), state space models (Mambas) (He et al., 2024a), diffusion models (Yin et al., 2023; He et al., 2024b), and other approaches (Guo et al., 2024; Zhao, 2023; He et al., 2024c). Nevertheless, the multi-class setting inevitably induces identity mapping in most methods, resulting in performance degradation. While previous studies (You et al., 2022; Guo et al., 2025) have endeavored to mitigate shortcut learning, the efficacy of their proposed techniques in complex and highly diverse scenarios remains limited (see Fig. 1 (c)-(d)). This is particularly evident in datasets such as Real-IAD and Universal Medical, where the abundance of normal patterns exacerbates the aforementioned issue (You et al., 2022). The effectiveness of simple noise operations (You et al., 2022; Guo et al., 2025) is diminished due to the enhanced noise robustness acquired through exposure to diverse visual patterns. Consequently, to enhance performance, our work aims to efficiently address the problem of identity shortcuts. Furthermore, why are the Real-IAD and Universal Medical datasets more complex? We elaborate on this as follows: It is apparent that in comparison to widely adopted industrial datasets—MVTec-AD (encompassing 15 classes) (Bergmann et al., 2019) and ViSA (comprising 12 classes) (Zou et al., 2022)—Real-IAD (Wang et al., 2024) exhibits more diverse normal patterns, featuring 30 object categories and 5 camera views. In contrast to standardized industrial images, although the Universal Medical dataset (He et al., 2024b) consists of only 3 categories, the inherent heterogeneity within medical normal samples enriches the visual features (e.g., variations in organ size and shape across different demographics, as illustrated in Fig. 2(a)). Moreover, the Universal Medical dataset displays a sparser t-SNE plot and a wider dispersion of feature distances compared to ViSA (see Fig. 2 (b) and (c)), indicating a higher degree of intrinsic diversity within medical data.
@@
In this paper, we propose a simple yet effective framework for the MUAD task, named ShortcutBreaker, built on the advanced DINO-pretrained vision transformer with two key innovations. First, drawing on two observed properties of low-rank matrix decomposition, we design a low-rank noisy bottleneck (LRNB) to effectively mitigate the identity mapping issue. Within LRNB, the low-rank property theoretically circumvents shortcut learning of unseen patterns, while learnable matrix parameters are optimized to reconstruct normal patterns. Second, we introduce a Frequency-Filter Unit (FFU) to replace self-attention, which efficiently curbs information leakage from input to output in the frequency domain. In FFU, a frequency spectral filtering and a local masking mechanism are integrated, forcing the decoder to learn reconstruction based on longer-range and impaired features.
To validate the effectiveness of the proposed method, we conduct extensive experiments on four publicly available datasets: MVTec-AD (Bergmann et al., 2019), ViSA (Zou et al., 2022), Universal Medical (He et al., 2024b), and Real-IAD (Wang et al., 2024). As presented in Fig. 1, our ShortcutBreaker achieves the highest image-level AUC: 99.8%, 98.9%, 88.2%, and 90.6% on these four datasets, respectively. Notably, on the complex Universal Medical and Real-IAD datasets, it outperforms previous methods by a significant margin. In summary, our contributions are:
@@
  • We propose a frequency-filtering unit that integrates frequency spectral filtering and global-local masking operations, detecting anomalies more efficiently than standard self-attention.
@@ 1.1 Multi Class UAD
UniAD (You et al., 2022) pioneers the field of multi-class unsupervised anomaly detection (MUAD), presenting a unified model capable of detecting anomalies across various classes. Most subsequent works have explored advanced modules to build better reconstruction models for MUAD. For example, LafitE (Yin et al., 2023) and DiAD (He et al., 2024b) further advance the MUAD task by leveraging the generative power of diffusion models to better capture anomalies across multiple classes. ViTAD (Zhang et al., 2023) and MambaAD (He et al., 2024a) develop feature reconstruction-based MUAD methods using recently advanced modules: Vision Transformer (Zhang et al., 2023) and State Space Model (He et al., 2024a), respectively. Few methods explicitly aim to address the identity mapping issue. UniAD (You et al., 2022) counteracts this through techniques like feature jittering and neighbor-masked attention. Dinomaly (Guo et al., 2025) employs components such as noisy bottleneck dropout to disrupt input feature replication. However, these noise operations have limited effectiveness in preventing shortcut learning when training on complex medical/industrial datasets (Universal Medical and Real-IAD).
@@ 2.1 Overview
Following ViTAD and Dinomaly (Zhang et al., 2023; Guo et al., 2025), our ShortcutBreaker is a feature-reconstruction framework and is constructed based on the vision transformer (ViT) structure. As depicted in Fig. 3(a), ShortcutBreaker consists of a pre-trained encoder, a bottleneck, and a decoder. The DINO-pretrained ViT model (Darcet et al., 2023) is utilized as the encoder, capturing informative feature tokens to facilitate subsequent reconstruction. The bottleneck is built upon a multi-layer perceptron (MLP), integrating multi-scale feature representations from intermediate encoder layers. The decoder based on frequency-filtering blocks maintains structural similarity with the encoder, employing transformer layers to reconstruct feature maps. As illustrated in Fig. 3(d), the encoder remains frozen during training, while the remaining components of ShortcutBreaker are optimized to reconstruct the encoder’s feature tokens by minimizing inter-token representation discrepancies. During inference, this configuration enables accurate reconstruction of normal patterns but exhibits reconstruction failures in anomalous regions. The anomaly localization map is generated by computing 1-CS (where CS denotes cosine similarity) between the original and reconstructed feature tokens and then reshaping the result. The anomaly detection score is subsequently determined by the maximum value within this anomaly map.
Figure 3: Flowchart of our ShortcutBreaker: (a) the overall structure of our proposed method; (b) the structure of the Low-Rank Noisy Bottleneck (LRNB); (c) the structure of the Frequency-Filtering Block (FFB), which consists of a Frequency-Filtering Unit and a Feedforward block. FFT: Fast Fourier Transform, iFFT: Inverse Fast Fourier Transform, FSF: frequency spectral filtering, LM: local masking $M_{l}$, and GM: global masking $M_{g}$.
@@ 2.2 Low-Rank Noisy Bottleneck (LRNB)
As discussed in the introduction, under the MUAD setting, identity shortcuts readily occur, narrowing the gap in anomaly scores between normal and abnormal cases. Previous methods (Guo et al., 2025; You et al., 2022) propose perturbing extracted feature tokens to compel the network to reconstruct from information-impaired normal features, rather than directly copying them. However, these noisy operations are often limited in effectively preventing shortcut learning in complex scenarios. We hypothesize that richer visual patterns inherent in the data enhance the reconstruction model’s robustness to stochastic noise. Consequently, we aim to design a novel paradigm for avoiding shortcuts that goes beyond solely relying on noisy operations. Through empirical observation of low-rank matrix decomposition in LoRA (Hu et al., 2022) (or the encoder-decoder structure in Auto-Encoder (He et al., 2022)), we identified two inherent properties particularly suitable for addressing the identity mapping issue.
Property 1 This property stems from the constrained low-rank latent space, which effectively circumvents identity shortcuts. In deep learning, nonlinear activations are generally used to improve performance, so we introduce the tool of the Jacobian matrix (Goodfellow et al., 2016) to conduct a linear approximation for the proof of this property. Given that the feature tokens extracted from DINO-v2 are highly informative, we assume them to be full-rank.
To achieve identity mapping of $x\in\mathbb{R}^{d}$, we require that $g\circ f(x)=x$ holds for all $x\in\mathbb{R}^{n}$. Differentiating both sides with respect to $x$ yields the Jacobian equation:
@@
$\tiny r\left(J_{g}(f(x))\cdot J_{f}(x)\right)\leq\min\left(r(J_{g}(f(x))),r(J_{f}(x))\right)\leq k$} (3)
@@
Here, $r(\cdot)$ indicate the rank of the input. Under the low-rank adaptation module, we have $k<d$, Substituting into Eq. (3) implies:
$\tiny r\left(J_{g}(f(x))\cdot J_{f}(x)\right)\leq k<d$} (4)
@@
This contradicts Eqs. (2) that $det(I_{d\times d})$ requires rank=$d$. Consequently, no such functions $f$ and $g$ exist.
@@
Property 2 Property 1 prevents the identity mapping of all the inputs; however, our final goal is a favorable score trade-off between normal and abnormal samples. The property 2 emerges from training exclusively on normal samples, during which the learnable $g$ and $f$ layers are optimized to specifically enhance the reconstruction of normal patterns. This optimization process minimizes adverse effects on the reconstruction fidelity of normal cases, thereby enlarging the gap of reconstruction errors between the two types of samples.
Specifically, functions $g$ and $f$ are parameterized by Multi-Layer Perceptrons (MLPs) with GeLU activations. As illustrated in Fig. 3, $g$ and $f$ feature symmetric architectures consisting of $L$ downsampling (DB) and upsampling blocks (UB), respectively. Within these modules, each DB reduces the token dims by low-rank ratio $r$, (i.e., $d_{i}=d_{i-1}//r$, sequence length $n$ remains unchanged), while UBs correspondingly scale it back. To reinforce feature-level perturbation, Dropout (Guo et al., 2025) is integrated into the UBs. Furthermore, departing from existing methods that employ a single noise bottleneck, we deploy dual LRNBs—injecting one before each of the two decoder groups—to perturb reconstruction.
2.3 Frequency-Filtering Block
The self-attention mechanism is the core component of Transformer, as shown in Eq. 5, enabling the model to focus on different parts of the input tokens.
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V=A(X)V$ (5)